# Perceptions of Linguistic Uncertainty by Language Models and Humans

### v1

@@ Abstract
Uncertainty expressions such as “probably” or “highly unlikely” are pervasive in human language. While prior work has established that there is population-level agreement in terms of how humans interpret these expressions, there has been little inquiry into the abilities of language models to interpret such expressions. In this paper, we investigate how language models map linguistic expressions of uncertainty to numerical responses. Our approach assesses whether language models can employ theory of mind in this setting: understanding the uncertainty of another agent about a particular statement, independently of the model’s own certainty about that statement. We evaluate both humans and 10 popular language models on a task created to assess these abilities. Unexpectedly, we find that 8 out of 10 models are able to map uncertainty expressions to probabilistic responses in a human-like manner. However, we observe systematically different behavior depending on whether a statement is actually true or false. This sensitivity indicates that language models are substantially more susceptible to bias based on their prior knowledge (as compared to humans). These findings raise important questions and have broad implications for human-AI alignment and AI-AI communication.
Keywords: Uncertainty Estimation, Large Language Models, Uncertainty Expression, Uncertainty Quantification, Theory of Mind, Human studies
@@ 1 Introduction
Uncertainty is ubiquitous in human communication — in relaying predictions (“it is likely to rain tomorrow”), conveying imperfect knowledge (“I think I have a copy in my desk”), and describing unknown information (“the artifact could be more than 500 years old”). Expressing uncertainty is critical in fields such as medicine, law, and politics, where statements including uncertainty expressions (e.g., “likely,” “doubtful”) are frequently used to support medical, judicial, and political decisions (Karelitz and Budescu, 2004). Domain experts use these expressions to communicate uncertainty across a variety of situations, such as the likelihood of side-effects of a medical treatment (Sawant and Sansgiry, 2018; Patt and Dessai, 2005), the chances of winning a not-guilty verdict in legal cases (Fore, 2019), the probability of environmental events resulting from climate change (Patt and Dessai, 2005; Ho et al., 2015), or the likelihood of emergence of military conflicts (Duke, 2023). Prior work has found that, in general, humans are well-attuned to the use of such uncertainty expressions, exhibiting population-level agreement in mapping these expressions to corresponding probabilities (Wallsten et al., 1986a; Willems et al., 2019; Fagen-Ulmschneider, 2019).
@@
However, the topic of how large language models (LLMs) interpret linguistic uncertainty has received relatively little attention (Sileo and Moens, 2023). In particular, given text where a speaker expresses uncertainty about a particular statement, this paper investigates whether LLMs can interpret the uncertainty not as a function of the model internal beliefs, but by objectively assessing the speaker’s uncertainty about the statement. Consider the motivating example in Figure 1: when writing a headline for a statement qualified by the word “probable,” ChatGPT expresses substantially different forms of linguistic uncertainty depending on its prior belief about the statement. [^1] When prompted about its belief about the different statements, ChatGPT agrees with the first one and disagrees with the second; see Figure 9 in the Appendix A. In this example, ChatGPT is conflating the speaker’s uncertainty with its own uncertainty about the statement—in effect, a failure of “theory of mind.”
@@
Figure 1: Two interactions with ChatGPT (June 2024). In each, ChatGPT is asked to write a headline for a short passage. Both passages are structured identically and qualified with the word “probable”, but the first is about climate change and the second about the link between vaccines and autism. For the first passage, ChatGPT generates a confident-sounding headline, using words like “conclude” and “comprehensive”. The second headline is weaker, with words like “suggests” and “possible”.
@@
In this work, we investigate the abilities of LLMs to provide quantitative interpretations of uncertainty expressions, focusing in particular on how the prior knowledge of an LLM affects this ability. To this end, we introduce a new task [^2] All code and data used in this study are available at https://github.com/UCIDataLab/llm-uncertainty-perceptions. in which LLMs must map text containing uncertainty expressions to numerical responses. We analyze the performance of both humans and 10 popular LLMs on this task, enabling direct comparison between humans and models. We find that larger, newer models like GPT-4 and Llama-3 can consistently map uncertainty expressions to numerical responses that align with human population-level perceptions. However, we also show that the responses LLMs generate are susceptible to bias based on their prior knowledge—to a much greater extent than those of humans.
@@
The models’ sensitivity to its prior knowledge has concerning implications given the increasing use of LLMs for generating content (e.g., summarization, data augmentation) (Wang et al., 2023; Shao et al., 2024; Laban et al., 2024), evaluating language generation (Zheng et al., 2023), simulating human behavior (Argyle et al., 2023; Park et al., 2023a), or even in building multi-agent pipelines (Wu et al., 2023; Qian et al., 2024). When an LLM’s ability to quantify uncertainty can be “poisoned” by its beliefs, its downstream performance is dependent on its parametric or pretraining knowledge (which can be obsolete or wrong Liang et al. (2022); Longpre et al. (2023)), rather than on critical contextual information (Longpre et al., 2021). Further, this means that the biases of a model (including the many well-documented potentially harmful biases of LLMs, e.g., Wan et al. (2023); Kotek et al. (2023); Salewski et al. (2024); Scherrer et al. (2024); Motoki et al. (2024)) can subtly manifest in how it interprets and generates uncertainty language and, as a consequence, have broader implications for human-AI and AI-AI interactions.
@@ 2 Related Work
In fields like medicine, finance, law, and politics, where it is impossible to make predictions with complete certainty, decisions are often informed by subjective probabilities (Karelitz and Budescu, 2004; Dhami and Wallsten, 2005; Fore, 2019). Subjective probabilities can be communicated quantitatively, through numerical probabilities (e.g., odds, percentages, intervals), or qualitatively, through the use of uncertainty expressions or epistemological markers (e.g., “I believe”, “According to”) (Dhami and Mandel, 2022).
Although being less precise than numerical probabilities (Wallsten et al., 1986b; Brun and Teigen, 1988; Budescu et al., 2014), humans generally prefer to use linguistic expressions, rather than numbers, to communicate uncertainty (Erev and Cohen, 1990; Wallsten et al., 1993). Interested in the efficacy of how humans communicate uncertainty linguistically, researchers have examined how individuals map uncertainty expressions into numerical values across different fields and expertise levels (Windschitl and Wells (1996); Karelitz and Budescu (2004); Wallsten et al. (2008, 1986a); Fore (2019); inter alia). Although there can be considerable variation in responses at the individual level, these studies have revealed that there are consistent patterns relating uncertainty expressions and numerical probabilities that can be observed systematically at the population level (Budescu and Wallsten, 1985; Clarke et al., 1992; Wallsten et al., 2008; Willems et al., 2019; Fagen-Ulmschneider, 2019).
@@
The need for more reliable LLMs has prompted researchers to investigate new methods for communicating the internal uncertainty of LLMs. Proposed methods differ in the information used to estimate the model’s uncertainty, ranging from methods that use token-level information (Jiang et al., 2021; Kuhn et al., 2023; Duan et al., 2024), to dissimilarities across multiple samples (Si et al., 2022; Chen and Mueller, 2023; Xiong et al., 2024; Hou et al., 2024; Lin et al., 2024; Aichberger et al., 2024), to training external classifiers using the inputs and/or LLMs’ representations (Jiang et al., 2021; Mielke et al., 2022; Shrivastava et al., 2023), or even directly eliciting confidence estimates from LLMs as output tokens (Lin et al., 2022; Tian et al., 2023). Furthermore, several works have analyzed the impact of LLM-articulated uncertainty in human-AI interaction, finding that participants adjust their perception of LLMs’ correctness when shown LLM outputs that include uncertainty expressions (Zhou et al., 2024; Kim et al., 2024; Steyvers et al., 2024). While these works investigate how LLMs express uncertainty when generating text and how humans react to various uncertainty expressions in text, there has been far less work on the questions we focus on in this paper, i.e., how LLMs interpret uncertainty in text and how closely these interpretations match those of humans.
@@
A small body of work has begun to investigate the relationships between uncertainty expressions and model behavior. Sileo and Moens (2023) investigate whether LLMs are able to discriminate between two uncertainty expressions and reason in terms of compositions of expressions. However, the paper focuses on LLMs’ binary rankings of expressions, as opposed to numerical interpretations, and does not compare LLM and human behavior. Most directly related to our work is that of Maloney et al. (2024) which compares numerical probability estimates from GPT-4 and humans using a small set of “context” prompts. Tang et al. (2024) also compare the numeric-textual mapping of uncertainty expressions across four different contexts in both Chinese and English languages. Our paper goes significantly beyond these works by assessing a broad range of LLMs using a more diverse and natural set of contexts. We additionally conduct human experiments, assessing the performance of humans on the same task to facilitate a direct comparison between humans and LLMs. Further, our approach is designed to target “theory of mind”—the task requires humans and LLMs to quantify what an uncertainty expression reflects about the speaker’s belief, rather than what the expression means to them. Finally, our work is the first that we are aware of to investigate how LLMs can be biased by their prior knowledge in mapping uncertainty expressions to numerical probabilities.
@@ 3 Baseline Human Study
As a baseline for how people map uncertainty expressions to numerical probabilities, we first conducted an experiment in which 94 humans were shown uncertainty expressions and asked to provide corresponding numerical responses. We focused on a set of 14 uncertainty expressions (e.g., “almost certain”, “unlikely”—the full list is provided in Appendix B and is also shown on the y-axis in Figure 4), drawn from Wallsten et al. (2008) and Wallsten et al. (1986a). In this initial experiment, our goal is to assess how people perceive these uncertainty expressions “in the wild,” putting them in the context of plausible real-world statements. In addition, we also aim to use statements that minimize the potential for people to conflate their own beliefs about these statements with their assessment of the confidence of the person making the statement.
@@
To this end, we constructed a set of statements $(u,s,e)$ which include uncertainty expressions $u\in\mathcal{U}$ used by speakers $s\in\mathcal{S}$ to convey their degree of certainty about the truthfulness or falsehood of a statement or event $e\in\mathcal{E}$. By presenting statements about a specific speaker $s$, [^3] Names are selected arbitrarily from a pre-defined pool of names. Participants see each name once throughout the experiment. For more details, see Appendix B. we are asking participants to use “theory of mind” to estimate how likely is that the speaker believes that the statement is true. We then query participants about the speaker’s strength of belief, clearly distinguishing this notion from the participant’s own beliefs. For instance, given the statement “Sonia believes it is unlikely it will rain today,” we can ask participants to quantify with a numerical response how likely Sonia thinks it is that it will rain today, distinct from the probability the participants themselves believe it will rain. We use the term numerical response throughout the paper to refer to the participant’s numerical assessment of the likelihood that speaker $s$ believes statement $e$ to be true when using uncertainty expression $u$. We instruct participants that responses should be numbers between 0 and 100, where 0 implies that a speaker $s$ believes there is a $0\%$ chance that the statement $e$ is true whilst 100 implies a speaker believes that there is a $100\%$ chance that the statement is true.
@@
In our baseline experiment, we use non-verifiable statements $e$ to separate the meaning of the uncertainty expressions from uncertainty about the statements themselves. These are statements that are not sufficiently grounded with specific contextual information to allow an external observer to be confident in either the truth or falsity of the statement. For example, in the context of a prompt such as “Maria believes it is likely that [statement]”, statements such as her boss has two pets or her flight will land around 6pm are statements we consider non-verifiable in the sense that there is insufficient context provided (that the speaker knows, but does not state) for an observer to be able to reliably assess the likelihood that the statement is true. In contrast, verifiable statements (which we discuss further in Section 4.1) can be verified as correct or incorrect in a context-free sense (e.g., the capital city of Peru is Lima); humans and LLMs will often have strong prior beliefs about the likelihood that such statements are true.
@@
Specifically, we manually constructed a set of 60 non-verifiable statements and systematically combined these with 14 uncertainty expressions. We randomly generated speaker names, generating sentences describing the belief of a hypothetical speaker in the form: “[Speaker] believes it is [uncertainty phrase] that [statement].” For each of these sentences, participants were asked to quantify the speaker’s belief about the statement, in particular, they were asked what is the probability being expressed from the speaker’s perspective that the statement is correct. Participants then provided their response quantized to numerical bins $0,5,10,\ldots,95,100$. Figure 2 illustrates one of the examples shown to participants in the experiment. Each of the 94 participants provided annotations for 28 randomly chosen statements (and speaker names). As a result, every uncertainty expression was annotated twice by each participant.
The outcome of this experiment [^4] Additional details about the experiments can be found in Appendix B. is an empirical distribution over the $2\times 94$ numerical responses that participants associated with each uncertainty expression. For example, Figure 3 reflects the histogram of responses assigned to the phrases “very likely” and “very unlikely”; results for all 14 uncertainty expressions can be summarized in a heat-map as shown in Figure 4. Overall, we observe results that are in agreement with prior work on human perceptions of uncertainty expressions (Wallsten et al., 2008, 1986a; Willems et al., 2019), including consistent ordering in aggregate population patterns, as determined by the mode of the empirical distributions.
@@
Figure 2: Example question of a non-verifiable statement provided to participants in the baseline experiment. Each example constitutes a unique name and statement. Participants see one question at a time.
(a) “very likely”
(b) “very unlikely”
Figure 3: Histogram of participant responses for non-verifiable statements for two uncertainty expressions.
@@
Figure 4: Human empirical distributions of numerical responses per uncertainty expression in the non-verifiable setting. Highlighted blue boxes represent the mode of each expression.

### v2

@@ Abstract
Uncertainty expressions such as “probably” or “highly unlikely” are pervasive in human language. While prior work has established that there is population-level agreement in terms of how humans quantitatively interpret these expressions, there has been little inquiry into the abilities of language models in the same context. In this paper, we investigate how language models map linguistic expressions of uncertainty to numerical responses. Our approach assesses whether language models can employ theory of mind in this setting: understanding the uncertainty of another agent about a particular statement, independently of the model’s own certainty about that statement. We find that 7 out of 10 models are able to map uncertainty expressions to probabilistic responses in a human-like manner. However, we observe systematically different behavior depending on whether a statement is actually true or false. This sensitivity indicates that language models are substantially more susceptible to bias based on their prior knowledge (as compared to humans). These findings raise important questions and have broad implications for human-AI and AI-AI communication.
Perceptions of Linguistic Uncertainty by Language Models and Humans
Catarina Belem*,1, Markelle Kelly*,1, Mark Steyvers1,2, Sameer Singh1, Padhraic Smyth1, 1Department of Computer Science, University of California Irvine 2Department of Cognitive Sciences, University of California Irvine
@@ 1 Introduction
The expression of uncertainty is ubiquitous in human communication — in relaying predictions (“it is likely to rain tomorrow”), conveying imperfect knowledge (“I think I have a copy in my desk”), and describing unknown information (“the artifact could be more than 500 years old”). 00footnotetext: Authors contributed equally to this work. Correspondence: cbelem@uci.edu. Expressing uncertainty is particularly critical in fields such as medicine, law, and politics, where statements including uncertainty expressions (e.g., “likely,” “doubtful”) are frequently used to support medical, judicial, and political decisions (Karelitz and Budescu, 2004). Domain experts use these expressions to communicate uncertainty across a variety of situations, such as the likelihood of side-effects of a medical treatment (Sawant and Sansgiry, 2018; Patt and Dessai, 2005), the chances of a not-guilty verdict in legal cases (Fore, 2019), the probability of environmental events resulting from climate change (Patt and Dessai, 2005; ho2015improving), or the likelihood of emergence of military conflicts (Duke, 2023). Prior work has found that, in general, humans are well-attuned to the use of such uncertainty expressions, exhibiting population-level agreement in mapping these expressions to corresponding probabilities (Wallsten et al., 1986a; Willems et al., 20; Fagen-Ulmschneider, 2019).
@@
However, the topic of how large language models (LLMs) interpret linguistic uncertainty has received relatively little attention. In particular, given text where a speaker expresses uncertainty about a particular statement, this paper investigates whether LLMs can interpret the uncertainty not as a function of the model’s internal beliefs, but by objectively assessing the speaker’s uncertainty about the statement. Consider the motivating example in Figure 1: when writing a headline for a statement qualified by the word “probable,” ChatGPT expresses substantially different uncertainty depending on its prior belief about the statement. [^1] When prompted about its belief about these statements, ChatGPT agrees with the first and disagrees with the second; see Figure 8 in the Appendix A. In this example, ChatGPT is conflating the speaker’s uncertainty with its own uncertainty about the statement—in effect, a failure of “theory of mind.”
@@
Figure 1: Two interactions with ChatGPT (June 2024) concerning the generation of a headline for a short passage. Both passages are structured identically and qualified with the word “probable,” but the first is about climate change and the second about the link between vaccines and autism. For the first passage, ChatGPT generates a confident-sounding headline, using the words “conclude” and “comprehensive.” The second headline is weaker, with words like “suggests” and “possible.”
@@
In this work, we investigate the abilities of LLMs to provide quantitative interpretations of uncertainty expressions, focusing in particular on how the prior knowledge of an LLM affects this ability. To this end, we propose to evaluate models’ capabilities as a function of their ability to map text containing uncertainty expressions to numerical responses. We analyze the performance of both humans and 10 popular LLMs on this task, enabling direct comparison between humans and models. [^2] The code and data can be found at https://github.com/UCIDataLab/llm-uncertainty-perceptions. We find that larger, newer models like GPT-4 and LLama3 (70B) can consistently map uncertainty expressions to numerical responses that align with human population-level perceptions. However, we also show that the responses LLMs generate are susceptible to bias based on their prior knowledge—with much greater susceptibility than that of humans.
@@
The models’ sensitivity to their prior knowledge has concerning implications for the use of LLMs in tasks in which they must process and generate text containing uncertainty expressions, e.g., summarizing scientific reports or writing news articles (Shao et al., 2024; Laban et al., 2024). When an LLM’s ability to quantify uncertainty can be “poisoned” by its beliefs, its downstream performance is dependent on its parametric or pretraining knowledge (which can be obsolete or wrong Liang et al. (2022); Longpre et al. (2023)), rather than on critical contextual information (Longpre et al., 2021). Further, this means that the biases of a model (including the many well-documented potentially harmful biases of LLMs, e.g., Wan et al. (2023); Kotek et al. (2023); Salewski et al. (2024); Scherrer et al. (2024); Motoki et al. (2024)) can subtly manifest in how it interprets and generates uncertainty language and, as a consequence, have broader implications for human-AI and AI-AI interactions.
@@ 2 Related Work
In fields like medicine, finance, law, and politics, where it is impossible to make predictions with complete certainty, decisions are often informed by subjective probabilities (Karelitz and Budescu, 2004; Dhami and Wallsten, 2005; Fore, 2019). Subjective probabilities can be communicated quantitatively, e.g., through numerical probabilities, odds, percentages, intervals, or qualitatively, through the use of uncertainty expressions or epistemological markers (e.g., “I believe”, “According to”) (Dhami and Mandel, 2022). Although they are less precise than numerical values (Wallsten et al., 1986b; Brun and Teigen, 1988; Budescu_Por_Broomell_Smithson_2014), humans generally prefer to use linguistic expressions to communicate uncertainty (Erev and Cohen, 1990; Wallsten et al., 1993).
Interested in the efficacy of how humans communicate uncertainty linguistically, researchers have examined how individuals map uncertainty expressions into numerical values across different fields and expertise levels (Windschitl and Wells (1996); Karelitz and Budescu (2004); Wallsten et al. (2008, 1986a); Fore (2019); inter alia). Although there can be considerable variation in responses at the individual level, these studies have revealed consistent and systematic patterns relating uncertainty expressions and numerical responses at the population level (Wallsten et al., 2008; Willems et al., 20; Fagen-Ulmschneider, 2019).
@@
The need for more reliable LLMs has prompted researchers to investigate new methods for communicating the internal uncertainty of LLMs. Proposed methods can be differentiated in terms of the information used to estimate the model’s uncertainty: from token-level information (Jiang et al., 2021; Kuhn et al., 2023; Duan et al., 2024), to dissimilarities across multiple samples (Si et al., 2022; Chen and Mueller, 2023; Xiong et al., 2024; Hou et al., 2024; Lin et al., 2024; Aichberger et al., 2024), to training external classifiers using the inputs and/or LLMs’ representations (Jiang et al., 2021; Mielke et al., 2022; Shrivastava et al., 2023), or even directly eliciting confidence estimates from LLMs as output tokens (Lin et al., 2022; Tian et al., 2023). Furthermore, several works have analyzed the impact of LLM-articulated uncertainty in human-AI interaction, finding that participants adjust their perception of LLMs’ correctness when shown LLM outputs that include uncertainty expressions (Zhou et al., 2024; Kim et al., 2024; Steyvers et al., 2024). With the goal of calibrating human-AI interaction,  Chaudhry et al. (2024) propose fine-tuning LLMs to convey uncertainty expressions that faithfully reflect their intrinsic uncertainty. While these works investigate how we can gauge LLMs’ intrinsic uncertainty and how humans react to various uncertainty expressions in text, there has been far less work on the questions we focus on in this paper, i.e., how LLMs interpret linguistic uncertainty and how closely these interpretations match those of humans.
@@
A small body of recent work has begun to investigate the relationships between uncertainty expressions and model behavior. Recently, Yona et al. (2024) found low correlations between LLMs’ intrinsic uncertainty and the use of uncertainty expressions. Sileo and Moens (2023) investigate whether LLMs are able to discriminate between two uncertainty expressions and reason in terms of compositions of expressions. However, the paper focuses on LLMs’ binary rankings of expressions, as opposed to numerical interpretations, and does not compare LLM and human behavior. Most directly related to our work is that of Maloney et al. (2024) which compares numerical probability estimates from GPT-4 and humans using a small set of “context” prompts, and the work of Tang et al. (2024) who compare the numeric-textual mapping of uncertainty expressions across four different contexts in both Chinese and English. Our paper goes significantly beyond this work by assessing a broad range of LLMs using a more diverse and natural set of contexts. We additionally conduct human experiments, assessing the performance of humans on the same task to facilitate a direct comparison between humans and LLMs. Further, our approach is designed to target “theory of mind”—the task requires humans and LLMs to quantify what an uncertainty expression reflects about the speaker’s belief, rather than what the expression means to the human or LLM. Finally, our work is the first that we are aware of to investigate how LLMs can be biased by their prior knowledge in mapping uncertainty expressions to numerical responses.
@@ 3 Baseline Human Study
As a baseline for how people map uncertainty expressions to numerical probabilities, we first conducted an experiment in which 94 humans were shown uncertainty expressions and asked to provide corresponding numerical responses. We focused on a set of 14 uncertainty expressions (e.g., “almost certain,” “unlikely”—the full list is provided in Appendix C and is also shown on the y-axis in Figure 3), drawn from Wallsten et al. (1986a) and Wallsten et al. (2008). In this initial experiment, our goal is to assess how people perceive these uncertainty expressions “in the wild,” putting them in the context of plausible real-world statements. In addition, we use types of statements that attempt to minimize the potential for people to conflate their own beliefs about these statements with their assessment of the confidence of the person making the statement.
@@
To this end, we constructed a set of statements $(u,s,e)$ which include uncertainty expressions $u\in\mathcal{U}$ used by speakers $s\in\mathcal{S}$ to convey their degree of certainty about the truthfulness or falsehood of a statement or event $e\in\mathcal{E}$. By presenting statements as being made by a specific speaker $s$, [^3] Names are selected arbitrarily from a pre-defined pool of names. Participants see each name once throughout the experiment. For more details, see Appendix C. we are asking participants to use theory of mind to estimate how likely it is that the speaker believes that the statement is true. We then query participants about the speaker’s degree of certainty, clearly distinguishing this notion from the participant’s own beliefs. For instance, given the statement “Sonia believes it is unlikely it will rain today,” we can ask participants to quantify with a numerical response how likely Sonia thinks it is that it will rain today, distinct from the participants’ own beliefs about how likely it is to rain. We use the term numerical response throughout the paper to refer to the participant’s numerical assessment of the likelihood that speaker $s$ believes statement $e$ to be true when using uncertainty expression $u$. We instruct participants that responses should be numbers between 0 and 100, where 0 implies that speaker $s$ believes there is a $0\%$ chance that statement $e$ is true while 100 implies the speaker believes there is a $100\%$ chance that the statement is true.
@@
In our baseline experiment, we use non-verifiable statements to separate the meaning of the uncertainty expressions from uncertainty about the statements themselves. Non-verifiable statements are statements that are not sufficiently grounded with specific contextual information to allow an external observer to be confident in either the truth or falsity of the statement. For example, in the context of a prompt such as “Maria believes it is likely that [statement],” we consider statements such as her boss has two pets or her flight will land around 6pm non-verifiable—as there is not enough context for an observer to be able to assess the likelihood that the statement is true. In contrast, verifiable statements (which we discuss further in Section 4.1) can be verified as correct or incorrect in a context-free sense (e.g., the capital city of Peru is Lima); humans and LLMs will often have strong prior beliefs about the likelihood that such statements are true.
@@
We manually constructed a set of 60 non-verifiable statements and systematically combined these with 14 uncertainty expressions. We randomly selected speaker names, generating sentences describing the belief of a hypothetical speaker in the form: “[Speaker] believes it is [uncertainty expression] that [statement].” For each sentence, participants were asked to quantify the speaker’s belief about the statement. In particular, they were asked what the probability is from the speaker’s perspective that the statement is correct. Participants then provided their response quantized to numerical bins $0,5,10,\ldots,95,100$ (see example in Figure 2). Each of the 94 participants provided annotations for 28 randomly chosen statements (and speaker names). As a result, every uncertainty expression was annotated twice by each participant.
@@
Figure 2: Example of a non-verifiable statement provided to participants in the baseline experiment. Each example uses a unique name and statement. Participants see one question at a time.
The outcome of this experiment [^4] Additional details about the experiments can be found in Appendix C. is an empirical distribution over the $2\times 94$ numerical responses that participants associated with each uncertainty expression. For example, Figure 9 reflects the histogram of responses assigned to the phrases “very likely” and “very unlikely”; results for all 14 uncertainty expressions can be summarized in a heat-map as shown in Figure 3. We note that our results are in agreement with prior work on human perceptions of uncertainty expressions (Wallsten et al., 2008, 1986a; Willems et al., 20), including consistent ordering in aggregate population patterns, in terms of the mode of the empirical distributions.
@@
Figure 3: Human empirical distributions of numerical responses per uncertainty expression in the non-verifiable setting. Highlighted blue boxes represent the mode value for each expression. Overall, population-level perceptions increase monotonically with the use of more confident uncertainty expressions.